OpenStudy (anonymous):

Could someone help me with this one question?!!!!? BEWLER? BEWLER? BEWLER? BEWLER?

OpenStudy (anonymous):



OpenStudy (anonymous):

use pythagoras theorem!

OpenStudy (anonymous):

is it cos= b/26 ?

OpenStudy (anonymous):

or sin= 20/26 ?

OpenStudy (anonymous):

b^2=26^2-20^2 b^2=676-400 now solve it yourself

OpenStudy (anonymous):

no i dont think you can use that because the angles have not been given.

OpenStudy (anonymous):

√b^2 = √276 b= 16.6133

OpenStudy (anonymous):

16.6 is the answer.

OpenStudy (anonymous):

Thank you